OpenStudy (anonymous):

se the quadratic formula to solve...2x^2-x=-2. I get halfway through and lose the answer it sems

OpenStudy (anonymous):

\[2x^2-x+2=0\] is a start, then use \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=2,b=-1,c=2\] unless you think this one factors easily (i don't)

OpenStudy (anonymous):

you get \[\frac{1\pm\sqrt{1-4\times 2\times 2}}{2\times 2}=\frac{1\pm\sqrt{-15}}{4}\]

OpenStudy (anonymous):

It doesn't, I spent forever on this

OpenStudy (anonymous):

As far as factoring easily

OpenStudy (anonymous):

which you can then write as \[\frac{1}{4}-\frac{\sqrt{15}}{4}i\]and its conjugate \[\frac{1}{4}+\frac{\sqrt{15}}{4}i\]